Assume the returns from holding small-company stocks are normally distributed. Also assume the average annual return for holding the small-company stocks for a period of time was 16.1 percent and the standard deviation of those stocks for the period was 34.6 percent. Use the NORMDIST function in Excel® to answer the following questions.


What is the approximate probability that your money will double in value in a single year? (Do not round intermediate calculations and enter your answer as a percent rounded to 3 decimal places, e.g., 32.161.)

Probability             %


What is the approximate probability that your money will triple in value in a single year? (Do not round intermediate calculations and enter your answer as a percent rounded to 8 decimal places, e.g., 32.16161616.)

Probability             %
Answer #1

Hello Sir/ Mam

Given that:

Mean = 16.10
Standard Deviation =                            34.60

(a)

P(x > 100) = 1–P(X < 100) = 1–P:5 )=1-P(75- 100 - 16.1 34.6 P(X > 100) = 1- P(Z < 2.424855) = 1 -0.9923 = 0.766%

(b)

P(x > 200) = 1-P(x < 200) = 1-P(75 )=1-P(2520 200 - 16.1 34.6 P(X > 200) = 1- P(Z < 5.315029) = 1-1=0 = 0%

I hope this solves your doubt.

Feel free to comment if you still have any query or need something else. I'll help asap.

Do give a thumbs up if you find this helpful.